OpenStudy (anonymous):

how would you write this quotient in standard form : 8-7i/1-2i

OpenStudy (anonymous):

\[\frac{ 8-7i }{ 1-2i }*\frac{ 1+2i }{ 1+2i }= \frac{ 8 +9i+14 }{ 1+4 }=\frac{ 22+9i }{ 5 }\]

OpenStudy (anonymous):

\[\frac{ 22 }{ 5 }+\frac{ 9i }{ 5 }\]

OpenStudy (anonymous):

where did the 9 come from?

OpenStudy (anonymous):

just simplified the multiplication in numerator

OpenStudy (anonymous):

16i-7i

OpenStudy (anonymous):

your sure that's right?

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

would the "I" be outside of the 9/5 ot connected to the 9 in the numerator

OpenStudy (anonymous):

connected to numerator

OpenStudy (anonymous):

okay thanks so much

OpenStudy (anonymous):

(;

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

i like your hair color, awesome ;)

OpenStudy (anonymous):

ahh thanks, I like to be different

OpenStudy (anonymous):

your different ;)

OpenStudy (anonymous):

how old are you?

OpenStudy (anonymous):

17 n u?

OpenStudy (anonymous):

the same!

OpenStudy (anonymous):

awesome (;

OpenStudy (anonymous):

quick question, if im finding the rational zeros and have already gotten far enough into the problem to where ive already done the synthetic division, and got the equation : 4x^4-4x^3-9x^2-1x+2, how would I factor that to get the zeros?

OpenStudy (anonymous):

your gonna get 4 zeroes here

OpenStudy (anonymous):

what would I factor?

OpenStudy (anonymous):

could I pull out an x?

OpenStudy (anonymous):

no u cant pull x out n it has got degree 4 so kinda complex

OpenStudy (anonymous):

so what do I do???

OpenStudy (anonymous):

helpppp

OpenStudy (anonymous):

no clue :(

OpenStudy (anonymous):

okay that's fine haha

OpenStudy (anonymous):

lol